calculate the mass of silver trioxonitrate (v) which reacts with 21.50g of sodium chloride to form 11.10g of silver chloride and 8.20g of sodium trioxonitrate (v) solution. Show that this reaction can be used to verify the law of conservation of matter
1
Expert's answer
2020-01-13T03:33:17-0500

AgNO3 + NaCl"\to" AgCl + NaNO3

m(AgNO3)=M(AgNO3)*m(AgCl)/M(AgCl)=169.87(g/mol)*11.1(g)/143.32(g/mol)=13.16g